Fermi National Laboratory


Accelerator Update - May 16-May 19

Last week Fermilab set a new integrated luminosity record with a level of 9159 nb-1.

Friday May 16
The day shift began with Tevatron (TeV) experts conducting end-of- store studies, with the Antiproton Source (Pbar) stacking, and with MiniBooNe taking beam.

At 10:12 AM, Operations terminated store 2549, and then at 10:53 AM put the TeV into shot setup.

Technicians began work on the Linac Klystron debuncher water system at 12:58 PM. MiniBooNe beam was off during this work. The work was completed at 1:20 PM.

At 1:45 PM, Operations established store 2551 with an initial luminosity of 43.08E30. Pbar resumed stacking.

Pbar sustained a stack-rate of 13.35 mA for one hour at 2:52 PM. This set a new stacking record.

At 3 PM, Recycler experts took shots of protons for studies.

Saturday May 17
The midnight shift began with Operations monitoring store 2551, with Pbar stacking, and with MiniBooNe taking beam.

At 7:33 AM, a Booster RF station (BRF13) tripped off several times due to a ground fault indication. Operators contacted experts.

Operations terminated store at 2551 at 8:05 AM.

At 8:22 AM, Operation had both the Collider Detector Facility (CDF) and D-zero (D0) collision halls ready for access.

Operations put the TeV through a dry squeeze at 9:12 AM. A TeV power supply (T:HE17) tripped off on a quench indication during the dry squeeze. Operators contacted experts.

At 10:31 AM, while an expert was investigating the problem with BRF13, it started working. Operations will keep an eye on the station.

Power supply techs put a diagnostics on T:HE17, and then Operations put the TeV through another dry squeeze at 12:56 PM. Operations then put the TeV into shot setup at 1:02 PM.

At 3 PM, Operations established store 2555 with an initial luminosity of 44.895E30. This set a new luminosity record.

Pbar set another new stacking-rate record with a rate of 13.51mA for one hour between 3:15 and 4:15 PM.

Sunday May 18
The midnight shift began with Operations monitoring store 2555, with Pbar stacking, and with MiniBooNe taking beam.

At 7:38 AM, a MI expert began conducting coalescing studies.

Operations terminated store 2555 and, at 8:10 AM, operators prepared the D0 collision hall for an access. The D0 experimenters were out of access by 9:20 AM.

Operations put the TeV through a dry squeeze at 9:26 AM, and then into shot setup at 10:06 AM.

Operators had to abort the TeV proton load when a sector D0 quadrupole (D0QT6) showed a fuse fault. Operators resolved the problem and continued with the shot setup.

At 12:36 PM, the TeV suffered an abort just after the last of the antiprotons were loaded due to a RF station (TRF3) trip; a power supply faulted. Pbar resumed stacking.

An expert began his investigation of the TRF3 fault. He didn't believe that TRF3 was the cause of the abort.

At 3:26 PM, Operations put the TeV through a dry squeeze.

A TeV expert arrived at 4:28 PM to begin orbit studies.

At 8:50 PM, the TeV suffered a sector BA quench while attempting to terminate the study beam.

A Linac RF station (LRF3) tripped off at 10:49 PM and then again at 11 PM. At 11:13 PM, LRF3 tripped off and would not reset. Operators called in a Linac tech.

Monday May 19
The midnight shift began with a Linac tech working on LRF3, with MiniBooNe taking beam, and with Pbar stacking.

The Linac tech had LRF3 back in operation by 1:35 AM.

Operations put the TeV into shot setup at 2:30 AM. Store 2562 was established at 3:56 AM. Its initial luminosity was 3o.07E30. Pbar resumed stacking.

At 4:11 AM, the Linac tech reported that LRF3 was still having problems. He said he'd come back during the morning and change out a driver tube.

LRF3 tripped off at 5:16 AM. Operators turned it and other stations off in preparation for the scheduled access day. They also began making MI ready for the Recycler access.

Today's Plan
The plan for today is to complete the complex accesses this evening and then began a five-shift study period at midnight. Store 2562 will remain in the TeV until the study period begins.

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/19/2003   email Fermilab